What capacitance when connected in series with a 500Ω resistor will limit the current

drawn from a 48-mV 465-kHz source to 20μA?



1
Expert's answer
2022-02-16T13:23:05-0500

Solution;

The impedance of the current is;

"Z=\frac VI =\sqrt{X_c^2+R^2}"

Substitute for the known values;

"Z=\frac{48×10^{-3}}{20×10^{-6}}=\sqrt{X_c^2+500^2}"

"2400=\sqrt{X_c^2+500^2}"

"X_c=\sqrt{2400^2-500^2}"

"X_c=2347.34\Omega"

But also;

"X_c=\frac{1}{2πfC}"

"C=\frac{1}{2πfX_c}=\frac{1}{2π×465000×2347.34}"

"C=1.458\u00d710^{-10}=0.1458nF"